Find all x ∈ R that solve this equation: 123 = x · (2x · (3x − 3) − 2) + 100 + 20 + 3


1
Expert's answer
2021-08-09T15:39:54-0400
"123 = x (2x (3x \u2212 3) \u2212 2) + 100 + 20 + 3"

"x (2x (3x \u2212 3) \u2212 2)=0"

"2x(3x^2-3x-1)=0"

"x=0\ or\ 3x^2-3x-1=0"

"x=\dfrac{3\pm\sqrt{(-3)^2-4(3)(-1)}}{2(3)}=\dfrac{3\pm\sqrt{21}}{6}"

"x\in \{0, \dfrac{3-\sqrt{21}}{6}, \dfrac{3+\sqrt{21}}{6}\}"